OpenStudy (anonymous):

A quadratic function and an exponential function are graphed below. Which graph most likely represents the quadratic function? f(x) because an increasing exponential function will eventually exceed an increasing quadratic function g(x) because an increasing quadratic function will eventually exceed an increasing exponential function g(x) because an increasing exponential f

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmmm I'm inclined towards the flatter one because an exponential function shoots up "y" values really fast while the quadratic is constrained to its "2" exponent or one can say for exampl, let's take x=8 so \(\large x=8\implies \begin{array}{llll} y=2^x\to 2^8\to 256 \\ \quad \\ y=x^2\to 8^2\to 64 \end{array}\)

OpenStudy (jdoe0001):

hmm I may have mixed up the values there lemme rewrite that

OpenStudy (jdoe0001):

so one can say the quadratic is g(x), since the quadratic is narrower, but that's only as they're close to 0

OpenStudy (anonymous):

So since the quadratic is g(x) would the answer be g(x) because an increasing quadratic function will eventually exceed an increasing exponential function g(x) because an increasing exponential function will always exceed an increasing quadratic function until their graphs intersect

OpenStudy (jdoe0001):

if you grab the graph, drag up up up and find where they intersect , which is at y=16 more or less you'll see the exponential gets narrower

OpenStudy (jdoe0001):

g(x) because an increasing exponential function \(\bf \text{will always exceed}\) an increasing quadratic function \(\bf \text{until their graphs intersect}\)

OpenStudy (anonymous):

ok thanks so the answer is g(x) because an increasing exponential function will always exceed an increasing quadratic function until their graphs intersect

OpenStudy (jdoe0001):

yes when close to 0, the quadratic will be narrower as they intersect further up, the exponential just takes off